# Splitting of uniform bundles on quadrics

### v1

@@ Abstract.
We show that there exist only constant morphisms from $\mathbb{Q}^{2n+1}(n\geq 1)$ to $\mathbb{G}(l,2n+1)$ if $l$ is even $(0<l<2n)$ and $(l,2n+1)$ is not $(2,5)$. As an application, we prove on $\mathbb{Q}^{2m+1}$ and $\mathbb{Q}^{2m+2}(m\geq 3)$, any uniform bundle of rank $2m$ splits.
1.Xinyi Fang is supported by innovation Action Plan (Basic research projects) of Science and Technology Commission of Shanghai Municipality (Grant No.21JC1401900)
@@
Keywords: uniform bundle; quadric; splitting of vector bundles.
@@ 1. Introduction
In this article, we assume all the varieties are defined over $\mathbb{C}$. Assume that $X$ is a rational homogeneous variety of Picard number $1$ (we call $X$ a generalised Grassmannian for short), then $X$ is swepted by lines. Let $E$ be a vector bundle on $X$. We consider its restriction $E|_{L}(\simeq\mathcal{O}(a_{1})\oplus\cdots\oplus\mathcal{O}(a_{k}))$ to any line $L\subseteq X$. If the splitting type $(a_{1},\dots,a_{k})$ of $E|_{L}$ is independent of the choice of $L$, we call $E$ a uniform bundle.
Every uniform bundle on $\mathbb{P}^{n}$ whose rank is smaller than $n$ splits (cf. the main theorems of [2] and [8]). On a Grassmannian $Gr(k,n+1)(k\leq n+1-k)$, any uniform bundle whose rank is smaller than $k$ splits (cf. [4, Theorem 1]). On some other generalised Grassmannians, there are similar results, see [1, Theorem 1.5], [5, Theorem 3.1] and [7, Theorem 5.6].
In a recent article [3], for an arbitrary generalised Grassmannian $X$, we prove that once the rank of a uniform bundle $E$ is smaller than $e.d.(\mathrm{VMRT})+1$ (for the definition of $e.d.(\mathrm{VMRT})$, see [3, Section $3$]), the bundle $E$ necessarily splits. Moreover, for most generalised Grassmannians, we classify unsplit uniform bundles of minimal ranks. The upper bound $e.d.(\mathrm{VMRT})$ is optimal for many generalised Grassmannians, for example it is optimal for $\mathbb{Q}^{n}(3\leq n\leq 6)$ (see [3, Theorem 1.1 (2)]). However, we will see that this upper bound is not optimal for higher dimensional quadrics.
In this article, we first show that there exist only constant morphisms from $\mathbb{Q}^{2n+1}(n\geq 1)$ to $\mathbb{G}(l,2n+1)$ if $l$ is even $(0<l<2n)$ and $(l,2n+1)$ is not $(2,5)$. Combining this result with the method of [2], we prove that every uniform bundle of rank $2n$ on $\mathbb{Q}^{2n+1}$ and $\mathbb{Q}^{2n+2}(n\geq 3)$ splits and the upper bound $2n$ is optimal for $\mathbb{Q}^{2n+1}$. This gives the first known example such that the optimal upper bound is bigger than $e.d.(\mathrm{VMRT})$.
This article is organised as follows: in Section $2,$ we study morphisms from $\mathbb{Q}^{2n+1}(n\geq 1)$ to $\mathbb{G}(l,2n+1)$ and the main result of this section is Proposition 2.1; in Section $3,$ we study uniform bundles on smooth quadrics and the main result of this section is Theorem 3.11. Furthermore, we also study uniform bundles on $B_{n}/P_{k}$ and $D_{n}/P_{k}$, see Corollary 3.13, Corollary 3.14, Corollary 3.15 and Corollary 3.16.
@@ 2. Morphisms from quadrics to Grassmannians
We mainly follow the proof of the main theorem in [10]. Let $H$ be the cohomology class of a hyperplane on $\mathbb{Q}^{2n+1}$. The cohomology ring of $\mathbb{Q}^{2n+1}$ is
@@
By [9, Proposition 2.1 (i)], both $c_{i}(1\leq i\leq l+1)$ and $d_{j}(1\leq j\leq 2n+1-l)$ are non-negative. We note that for any $1\leq i,j<n$, $c_{i}$ and $d_{j}$ are integers. For any $i,j\geq n$, $2c_{i}$ and $2d_{j}$ are integers. Then from the exact sequence $0\rightarrow f^{*}U\rightarrow\mathcal{O}_{\mathbb{Q}^{2n+1}}^{\oplus 2n+2}\rightarrow f^{*}Q\rightarrow 0$, we get the equality of polynomials:
@@
If $c_{l+1}d_{2n+1-l}$ is 00, then by (2.1) both $c_{1}$ and $d_{1}$ are zero, which implies that $f$ is constant. So we may assume the numbers $c_{1},d_{1},c_{l+1}$ and $d_{2n+1-l}$ are non-zero. By [9, Proposition 2.1 (ii)], all the rational numbers $c_{1},c_{2},\dots,c_{l+1}$ and $d_{1},d_{2},\dots,d_{2n+1-l}$ are positive.
Let $a$ be $2n+2]{c_{l+1}d_{2n+1-l}}$. We set $C_{i}:=\frac{c_{i}}{a^{i}}(1\leq i\leq l+1)$ and $D_{j}:=\frac{d_{j}}{a^{j}}(1\leq j\leq 2n+1-l)$. We note that if $a,C_{i}$ and $D_{j}$ are all positive integers, then by the same proof of [10, Case (ii)], we can get a contradiction. Then it suffices to show that $a,C_{i}$ and $D_{j}$ are integers.
We firstly show that $a$ is an integer. Since $l$ is even, similar to the proof in [9, Lemma 3.3 (ii) and Case 1 of (iii)], we can show that $a$ is $\frac{c_{m+1}}{c_{m}}$, where $m$ is $\frac{l}{2}$. So $a$ is rational, we may assume $a=s/t$, where $s$ and $t$ are coprime positive integers. By definition, we have $a^{2n+2}=c_{l+1}d_{2n+1-l}$. Since $2c_{l+1}$ and $2d_{2n+1-l}$ are integers, $4a^{2n+2}$ is an integer. So $t^{2n+2}$ divides $4$, which implies that $t$ is $1$, as $n$ is at least $1$. Hence $a$ is an integer.
We now show that $C_{i}(=\frac{c_{i}}{a^{i}})$ and $D_{j}(=\frac{d_{j}}{a^{j}})$ are integers. Let $F_{1}(x)\cdots F_{k}(x)$ be the irreducible factorization of $1-x^{2n+2}$ over $\mathbb{Z}[x]$ with $F_{l}(0)=1(1\leq l\leq k)$. Then $F_{l}(x)$ is also irreducible over $\mathbb{Q}[x]$ by Gauss Lemma. Then
@@ 3. Uniform bundles on $\mathbb{Q}^{2n+1}(n\geq 3)$
We are going to use the method in [2] to show that any uniform bundle of rank $2n$ on $\mathbb{Q}^{2n+1}(n\geq 3)$ splits. We firstly fix some notations.
Let $E$ be a uniform bundle on $X(=\mathbb{Q}^{2n+1})(n\geq 3)$ of rank $2n$. Assume that $E$ does not split, by [3, Proposition $3.7$], we may assume that the splitting type of $E$ is
@@
By [3, Lemma $2.1$], we have the following description of cohomology rings:
@@
If $a$ is 00, then both $S_{1}(T,X_{1},X_{2})$ and $S_{2}(T+X_{1},X_{1},X_{2})$ are polynomials only in variables $T$ and $X_{1}$. Since $S_{i}(T,X_{1},X_{2})(i=1,2)$ are symmetrical in $X_{1}$ and $X_{2}$, we must have $S_{1}(T,X_{1},X_{2})=T^{l+1}$ and $S_{2}(T,X_{1},X_{2})=T^{2n-l-1}$. So $c_{1}(E_{1})$ and $c_{1}(E_{2})$ are 00, and $\psi_{x}$ is constant for each $x\in X$, which implies that $E$ splits (see, for example, [3, Proposition $3.5$]). Therefore, we have $a\neq 0$.
@@ 3.1. Approximate solutions
A non-zero homogeneous polynomial $P(T,X_{1})$ with rational coefficients in variables $T$ and $X_{1}$ of degree $2n$ is called an approximate solution if $P(T,X_{1})-R(X_{1},X_{2})$ has a proper divisor $S(T,X_{1},X_{2})$ which is symmetrical in $X_{1}$ and $X_{2}$. We call such a divisor a symmetrical divisor.
@@
By the the equation (*3), both $\frac{1}{a}E(T,X_{1})$ and $\frac{1}{a}E(T-X_{1},X_{1})$ are approximate solutions. We call them approximate solutions associated with (*3).
@@
Let $P(T,X_{1})=\sum_{k=0}^{2n}p_{k}X_{1}^{k}T^{2n-k}$ be an approximate solution. Then one of the followings holds.
$(1)$ Any symmetrical divisor of $P(T,X_{1})-R(X_{1},X_{2})$ is of degree one.
@@
Let $S(T,X_{1},X_{2})$ be a symmetrical divisor of $P(T,X_{1})-R(X_{1},X_{2})$. Then $S(0,X_{1},X_{2})$ divides $p_{2n}X_{1}^{2n}-R(X_{1},X_{2})$. As $S(T,X_{1},X_{2})$ is symmetrical in $X_{1}$ and $X_{2}$, $S(0,X_{1},X_{2})$ divides $p_{2n}X_{2}^{2n}-R(X_{1},X_{2}).$ Therefore $S(0,X_{1},X_{2})$ divides $p_{2n}(X_{1}^{2n}-X_{2}^{2n})$. By the equation $R(X_{1},X_{2})(X_{1}^{2}-X_{2}^{2})=X_{1}^{2n+2}-X_{2}^{2n+2}$, we have
@@
If $p_{2n}$ is not 00 and $1$, we have $S(0,X_{1},X_{2})\mid X_{1}^{2n}-X_{2}^{2n}$ and $S(0,X_{1},X_{2})\mid X_{1}^{2n}(X_{1}-X_{2})(X_{1}+X_{2})$. Since $S(0,X_{1},X_{2})$ is symmetrical in $X_{1}$ and $X_{2}$, $S(0,X_{1},X_{2})$ is $c(X_{1}+X_{2})$ for some $c\in\mathbb{Q}$. In particular, $\deg(S)$ is $1$.
@@
We call an approximate solution $P(T,X_{1})=\sum_{k=0}^{2n}p_{k}X_{1}^{k}T^{2n-k}$ a primitive approximate solution if $p_{2n}\in\{0,1\}$ and $P(T,X_{1})-R(X_{1},X_{2})$ has a symmetrical divisor $S_{0}(T,X_{1},X_{2})$ such that there is a $2(n-p_{2n}+1)$-th primitive unit root $y_{0}$ satisfying $S_{0}(0,1,y_{0})=0$.
@@
The proof of Proposition 3.5 is similar to that of [2, Proposition 6.1 and Proposition 6.2], we leave them in the Appendix (see Propositions A.1 and A.2).

### v2

# uniform bundles on quadrics

@@ Abstract.
We show that there exist only constant morphisms from $\mathbb{Q}^{2n+1}(n\geq 1)$ to $\mathbb{G}(l,2n+1)$ if $l$ is even $(0<l<2n)$ and $(l,2n+1)$ is not $(2,5)$. As an application, we prove on $\mathbb{Q}^{2m+1}$ and $\mathbb{Q}^{2m+2}(m\geq 3)$, any uniform bundle of rank at most $2m$ splits, which improves the upper bound of splitting for uniform bundles obtained by Kachi and Sato [7]. We classify all unsplit uniform bundles of minimal rank on $B_{n}/P_{k}$ $(k=\frac{2n}{3},k\geq 6)$ and $D_{n}/P_{k}$ $(k=\frac{2n-2}{3},k\geq 6)$. We partially answer a conjecture of Ellia, which predicts that some uniform bundles of special splitting types on $\mathbb{P}^{n}$ necessarily split and we find some restrictions on the splitting types of unsplit uniform bundles of minimal rank.
1.Xinyi Fang is supported by National Natural Science Foundation of China (Grant No. 12471040)
@@
Keywords: uniform bundle; quadric; generalized Grassmannian.
@@ 1. Introduction
In this article, we assume all the varieties are defined over $\mathbb{C}$. We assume that $X$ is a rational homogeneous variety of Picard number $1$ and we call $X$ a generalized Grassmannian for short. It is well known that $X$ is swept by lines. Let $E$ be a vector bundle on $X$. We consider its restriction $E|_{L}(\simeq\mathcal{O}_{L}(a_{1}(L)\oplus\cdots\oplus\mathcal{O}_{L}(a_{r}(L)))$ to any line $L\subseteq X$. If the splitting type $(a_{1}(L),\dots,a_{r}(L))$ of $E|_{L}$ is independent of the choice of $L$, $E$ is called a uniform bundle. We say a vector bundle splits if it can be decomposed as a direct sum of line bundles. We say a vector bundle does not split or a vector bundle is unsplit if it cannot be decomposed as a direct sum of line bundles.
In [5], Grothendieck shows that every vector bundle on a projective line splits. However, for higher dimensional projective spaces, the splitting of vector bundles is far more intricate. In [15], Van de Ven studies the splitting property of uniform bundles. He demonstrates that every uniform bundle of rank $2$ on $\mathbb{P}^{n}$ $(n\geq 3)$ splits and every uniform bundle of rank $2$ on $\mathbb{P}^{2}$ is isomorphic to $\mathcal{O}_{\mathbb{P}^{2}}(a)\oplus\mathcal{O}_{\mathbb{P}^{2}}(b)$ or $T_{\mathbb{P}^{2}}(a)$ for some integers $a$ and $b$. The subsequent work by Sato [12] and Elencwajg, Hirschowitz, Schneider [2] extends these results: every uniform bundle on $\mathbb{P}^{n}$ of rank smaller than $n$ splits, while every unsplit uniform bundle on $\mathbb{P}^{n}$ of rank $n$ is isomorphic to $T_{\mathbb{P}^{n}}(a)$ or $\Omega_{\mathbb{P}^{n}}(b)$ for some integers $a$ and $b$.
Motivated by these advances, we address two central problems for uniform bundles on a generalized Grassmannian $X$.
  • Problem $1$: Determine the splitting threshold $\mu(X)$ such that any uniform bundle of rank at most $\mu(X)$ splits and there exists an unsplit uniform vector bundle of rank $\mu(X)+1$.
  • Problem $2$: Classify all uniform bundles of rank $\mu(X)+1$.
Let $\mathbb{G}(k-1,n-1)$ be the Grassmannian of $k$-dimensional subspaces of $\mathbb{C}^{n}$. In [6], Guyot solves the above problems for Grassmannians $X=\mathbb{G}(k-1,n-1)(k\leq n-k)$. For orthogonal Grassmannians $OG(n,2n+1)$, Muñoz, Occhetta and Solá Conde [8] achieve similar results.
Let $\mathbb{Q}^{n}$ be a projective smooth quadric of dimension $n$, with $n\geq 5$ odd (resp. even). Kachi and Sato [7, Theorem 4.1] prove that any uniform vector bundle on $\mathbb{Q}^{n}$ of rank at most $n-2$ (resp. $n-3$) splits. Muñoz, Occhetta and Solá Conde provide an alternative proof via a general splitting criterion for low-rank uniform bundles on varieties covered by lines (see [8, Corollary 3.3]). However we will demonstrate their upper bounds are not optimal. Actually, we show that for $\mathbb{Q}^{2n+1}$ and $\mathbb{Q}^{2n+2}$ $(n\geq 3)$, every uniform bundle of rank $2n$ splits.
In our recent work [4], we systematically address Problem $1$ and Problem $2$ for all generalized Grassmannians. A key insight is establishing a connection between the geometry of VMRT (variety of minimal rational tangents) and the splitting behavior of uniform bundles. Specifically, we demonstrate that for an arbitrary generalized Grassmannian $X$, a uniform bundle $E$ of rank at most the so-called $e.d.(\mathrm{VMRT})$ (for the definition of $e.d.(\mathrm{VMRT})$, see [4, Section $3$]) must split. Furthermore, for the majority of generalized Grassmannians, the bound $e.d.(\mathrm{VMRT})$ coincides with $\mu(X)$. Our approach diverges from Guyot’s algebraic strategy by adopting a more geometric and direct methodology. Notably, for most generalized Grassmannians whose VMRTs are products of several irreducible varieties, we reduce Problem $2$ to the classification of uniform bundles on projective spaces and quadrics, see [4, Proposition $4.4$]. This reduction highlights the pivotal role of quadrics in resolving the problem.
Given a generalized Grassmannian $X$, assume $\mu(X)$ is $e.d.(\text{VMRT})$. Let $E$ be an unsplit uniform bundle of minimal rank on $X$, that is, the rank of $E$ is $\mu(X)+1$. In [4, Proposition 3.8], we show that, up to twisting by a suitable line bundle, the splitting type of $E$ is $(1,\dots,1,0,\dots,0)$. Inspired by this result, when $\mu(X)$ is not necessarily $e.d.(\text{VMRT})$, we show that there are some restrictions on the splitting type of $E$. To be concrete, if the splitting type of $E$ is $(a_{1},a_{2},\dots,a_{r})$ $(a_{1}\geq a_{2}\geq\cdots\geq a_{r})$, then $max\{a_{i}-a_{i+1}|1\leq i\leq r-1\}$ is $1$. By using the same strategy, we give an affirmative partial answer to a conjecture of Ellia (see [3, Page 29, Conjecture]), which predicts that every uniform bundle on $\mathbb{P}^{n}$ of splitting type $(\underbrace{a_{1},\dots,a_{1}}_{l_{1}},\underbrace{a_{2},\dots,a_{2}}_{l_{2}},\dots,\underbrace{a_{k},\dots,a_{k}}_{l_{k}})$ with $a_{i}>a_{i+1}$ and $l_{i}\leq n-1$ for any $1\leq i\leq k$ necessarily splits. Motivated by Ellia’s conjecture, we propose a similar conjecture for any generalized Grassmannian as follows:
Conjecture: For an arbitrary generalized Grassmannian $X$, there exists a maximal positive integer $\nu(X)$ such that every uniform bundle of splitting type
$(\underbrace{a_{1},\dots,a_{1}}_{l_{1}},\underbrace{a_{2},\dots,a_{2}}_{l_{2}},\dots,\underbrace{a_{k},\dots,a_{k}}_{l_{k}})$
with $a_{i}>a_{i+1}$ and $l_{i}\leq\nu(X)$ for any $1\leq i\leq k$ necessarily splits.
In this article, we establish:
Main Results:
  1. (1)
    For $\mathbb{Q}^{2n+1}$ and $\mathbb{Q}^{2n+2}$ $(n\geq 3)$, every uniform bundle of rank $2n$ splits, with $\mu(\mathbb{Q}^{2n+1})=2n$.
  2. (2)
    For $B_{n}/P_{k}$ $(k=\frac{2n}{3})$ and $D_{n}/P_{k}$$(k=\frac{2n-2}{3})$, we classify all unsplit uniform bundles of minimal rank.
  3. (3)
    Give an affirmative partial answer to a conjecture of Ellia in a more general setting and find some restrictions on the splitting types of unsplit uniform bundles of minimal rank.
Note that for $\mathbb{Q}^{2n+1}$, $e.d.(\mathrm{VMRT})$ is $12n-$ 1 and the number $\mu(\mathbb{Q}^{2n+1})$ is $2n$. So this is the first known example such that the optimal upper bound $\mu(X)$ is bigger than $e.d.(\mathrm{VMRT})$.
This article is structured as follows:
  1. (1)
    Section $2$: Analyze morphisms from $\mathbb{Q}^{2n+1}$ to $\mathbb{G}(l,2n+1)$ and the main result of this section is Proposition 2.1.
  2. (2)
    Theorem 3.11: Prove splitting theorems for quadrics via relative Harder-Narasimhan filtrations and approximate solutions.
  3. (3)
    Corollaries 3.133.163.133.163.13–3.163.13 – 3.16: Extend classifications to generalized Grassmannians $B_{n}/P_{k}$ and $D_{n}/P_{k}$.
  4. (4)
    Section $4$: Prove some uniform bundles of special splitting types necessarily split.
@@ 2. Morphisms from quadrics to Grassmannians
We follow the proof of the main theorem in [14]. Let $H$ denote the cohomology class of a hyperplane on $\mathbb{Q}^{2n+1}$. The cohomology ring of $\mathbb{Q}^{2n+1}$ is
@@
By [13, Proposition 2.1 (i)], all coefficients $c_{i}(1\leq i\leq l+1)$ and $d_{j}(1\leq j\leq 2n+1-l)$ are non-negative. We note that for any $1\leq i,j<n$, $c_{i}$ and $d_{j}$ are integers. For any $i,j\geq n$, $2c_{i}$ and $2d_{j}$ are integers. Then from the exact sequence $0\rightarrow f^{*}U\rightarrow\mathcal{O}_{\mathbb{Q}^{2n+1}}^{\oplus 2n+2}\rightarrow f^{*}Q\rightarrow 0$, we get the equality of polynomials:
@@
If $c_{l+1}d_{2n+1-l}$ is 00, then from (2.1), we obtain both $c_{1}$ and $d_{1}$ are zero by induction, which implies that $f$ is constant. So we may assume the numbers $c_{1},d_{1},c_{l+1}$ and $d_{2n+1-l}$ are non-zero. By [13, Proposition 2.1 (ii)], all the rational numbers $c_{1},c_{2},\dots,c_{l+1}$ and $d_{1},d_{2},\dots,d_{2n+1-l}$ are positive.
Let $a$ be $2n+2]{c_{l+1}d_{2n+1-l}}$. We set $C_{i}:=\frac{c_{i}}{a^{i}}(1\leq i\leq l+1)$ and $D_{j}:=\frac{d_{j}}{a^{j}}(1\leq j\leq 2n+1-l)$. We note that if $a,C_{i}$ and $D_{j}$ are all positive integers, then by the same proof of [14, Page 204, Case (ii)], we can get a contradiction (for details, see Appendix A, Claim). Then it suffices to show that $a,C_{i}$ and $D_{j}$ are integers.
We first show that $a$ is an integer. Since $l$ is even, similar to the proof in [13, Lemma 3.3 (ii) and Case 1 of (iii)], we can show that $a$ is $\frac{c_{m+1}}{c_{m}}$ (see Appendix A, Corollary A.2), where $m$ is $\frac{l}{2}$. So $a$ is rational, we may assume $a=s/t$, where $s$ and $t$ are coprime positive integers. By definition, we have $a^{2n+2}=c_{l+1}d_{2n+1-l}$. Since $2c_{l+1}$ and $2d_{2n+1-l}$ are integers, $4a^{2n+2}$ is an integer. So $t^{2n+2}$ divides $4$, which implies that $t$ is $1$, as $n$ is at least $1$. Hence $a$ is an integer.
We now show that $C_{i}(=\frac{c_{i}}{a^{i}})$ and $D_{j}(=\frac{d_{j}}{a^{j}})$ are integers. Let $F_{1}(x)\cdots F_{k}(x)$ be the irreducible factorization of $1-x^{2n+2}$ over $\mathbb{Z}[x]$ with $F_{l}(0)=1(1\leq l\leq k)$. Then $F_{l}(x)$ is also irreducible over $\mathbb{Q}[x]$ by Gauss’s Lemma. Then
@@ 3. Uniform bundles on $\mathbb{Q}^{2n+1}(n\geq 3)$
We are going to use the method in [2] to show that any uniform bundle of rank $2n$ on $\mathbb{Q}^{2n+1}(n\geq 3)$ splits. We first fix some notations.
Let $E$ be a uniform bundle on $X(=\mathbb{Q}^{2n+1})(n\geq 3)$ of rank $2n$. Assume that $E$ does not split, by [4, Proposition $3.7$], we may assume that the splitting type of $E$ is
@@
By [4, Lemma $2.1$], we have the following description of cohomology rings:
@@
If $a$ is 00, then both $S_{1}(T,X_{1},X_{2})$ and $S_{2}(T+X_{1},X_{1},X_{2})$ are polynomials only in variables $T$ and $X_{1}$. Since $S_{i}(T,X_{1},X_{2})(i=1,2)$ are symmetric in $X_{1}$ and $X_{2}$, we must have $S_{1}(T,X_{1},X_{2})=T^{l+1}$ and $S_{2}(T,X_{1},X_{2})=T^{2n-l-1}$. So $c_{1}(E_{1})$ and $c_{1}(E_{2})$ are 00, and $\psi_{x}$ is constant for each $x\in X$, which implies that $E$ splits (see, for example, [4, Proposition $3.5$]). Therefore, we have $a\neq 0$.
@@ 3.1. Approximate solutions
A non-zero homogeneous polynomial $P(T,X_{1})$ with rational coefficients in variables $T$ and $X_{1}$ of degree $2n$ is called an approximate solution if $P(T,X_{1})-R(X_{1},X_{2})$ has a proper divisor $S(T,X_{1},X_{2})$ which is symmetric in $X_{1}$ and $X_{2}$. We call such a divisor a symmetric divisor.
@@
By the equation (*3), both $\frac{1}{a}E(T,X_{1})$ and $\frac{1}{a}E(T-X_{1},X_{1})$ are approximate solutions. We call them approximate solutions associated with (*3).
@@
Let $P(T,X_{1})=\sum_{k=0}^{2n}p_{k}X_{1}^{k}T^{2n-k}$ be an approximate solution. Then one of the following holds.
$(1)$ Any symmetric divisor of $P(T,X_{1})-R(X_{1},X_{2})$ is of degree one.
@@
Let $S(T,X_{1},X_{2})$ be a symmetric divisor of $P(T,X_{1})-R(X_{1},X_{2})$. Then $S(0,X_{1},X_{2})$ divides $p_{2n}X_{1}^{2n}-R(X_{1},X_{2})$. As $S(T,X_{1},X_{2})$ is symmetric in $X_{1}$ and $X_{2}$, $S(0,X_{1},X_{2})$ divides $p_{2n}X_{2}^{2n}-R(X_{1},X_{2}).$ Therefore $S(0,X_{1},X_{2})$ divides $p_{2n}(X_{1}^{2n}-X_{2}^{2n})$. By the equation $R(X_{1},X_{2})(X_{1}^{2}-X_{2}^{2})=X_{1}^{2n+2}-X_{2}^{2n+2}$, we have
@@
If $p_{2n}$ is neither 00 nor $1$, we have $S(0,X_{1},X_{2})\mid X_{1}^{2n}-X_{2}^{2n}$ and $S(0,X_{1},X_{2})\mid X_{1}^{2n}(X_{1}-X_{2})(X_{1}+X_{2})$. Since $S(0,X_{1},X_{2})$ is symmetric in $X_{1}$ and $X_{2}$, $S(0,X_{1},X_{2})$ is $c(X_{1}+X_{2})$ for some $c\in\mathbb{Q}$. In particular, $\deg(S)$ is $1$.
@@
We call an approximate solution $P(T,X_{1})=\sum_{k=0}^{2n}p_{k}X_{1}^{k}T^{2n-k}$ a primitive approximate solution if $p_{2n}\in\{0,1\}$ and $P(T,X_{1})-R(X_{1},X_{2})$ has a symmetric divisor $S_{0}(T,X_{1},X_{2})$ such that there is a $2(n-p_{2n}+1)$-th primitive unit root $y_{0}$ satisfying $S_{0}(0,1,y_{0})=0$.
@@
The proof of Proposition 3.5 is similar to that of [2, Proposition 6.1 and Proposition 6.2], we leave them in the Appendix (see Propositions B.1 and B.2).